OpenStudy (anonymous):

Write an equation that would be solved by using multiplication. Then, show how to solve the equation.

OpenStudy (anonymous):

x/5=4 5x4=20 20/5=4

OpenStudy (salazarblack):

any help? http://openstudy.com/updates/53b848ebe4b09f140ca63c4f

OpenStudy (anonymous):

33.04 divided by 11.8 = b b= \[2.8\]

OpenStudy (anonymous):

thanks eveyone

OpenStudy (anonymous):

any more questions?

OpenStudy (anonymous):

\[11.8 \times 2.8= 33.08\]

OpenStudy (anonymous):

um no not for math anyway

OpenStudy (salazarblack):

i do but apparently no-one 's being helpful :/

OpenStudy (anonymous):

oops i meant 33.04